# The Jaynes-Cummings model as an autonomous Maxwell demon

### v2

@@ Abstract
$\hat{\rho}_{C}\left(t\right)$ $=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)+$ (S11)
$\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
@@
$\langle a\rangle$ $=Tr[\hat{\rho}_{C}\left(t\right)a]$ (S12)
$=\sum_{n}[\cos\left(gt\sqrt{n+1}\right)\cos\left(gt\sqrt{n}\right)+{\sin\left({gt\sqrt{n+1}}\right)}\sin\left({gt\sqrt{n}}\right)\frac{\sqrt{n+1}}{\sqrt{n}}]\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle$
@@
$\langle a\rangle$ $=\sum_{n}\left(\Lambda_{0}+\Lambda_{1}n+\Lambda_{2}n^{2}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle+O\left(\frac{1}{n_{0}^{4}}\right)$ (S13)
@@
$\Lambda_{0}=\frac{{1+\frac{\theta^{2}}{4}+2n_{0}\left(-3+4n_{0}\left(3+4n_{0}\right)-\theta^{2}\right)+\left(-1+6n_{0}-24n_{0}^{2}+\left(1-2n_{0}\right)^{2}\frac{\theta^{2}}{4}\right)\cos\left(\theta\right)-\left(1-3n_{0}\right)^{2}\theta\sin\left(\theta\right)}}{{32n_{0}^{3}}}$ (S14)
$\Lambda_{1}=\frac{{1-6n_{0}+\frac{\theta^{2}}{4}+\left(-1+\frac{\theta^{2}}{4}-2n_{0}\left(-3+\frac{\theta^{2}}{4}\right)\right)\cos\left(\theta\right)+\left(-2+7n_{0}\right)\frac{\theta}{2}\sin\left(\theta\right)}}{{8n_{0}^{3}}}$ (S15)
$\Lambda_{2}=\frac{{4+2\left(\frac{\theta^{2}}{4}-2\right)\cos\left(\theta\right)-\frac{5\theta}{2}\sin\left(\theta\right)}}{{16n_{0}^{3}}}$ (S16)
@@
$\langle a\rangle$ $=\sqrt{n_{0}}\Big{(}1+\frac{\sin^{2}{\frac{\theta}{2}}}{2{n_{0}}}+\frac{-2\sin^{2}{\frac{\theta}{2}}\left(1+\frac{\theta^{2}}{2}\right)+\frac{\theta}{2}\sin{\theta}+\bar{n}\theta\left(\theta\cos{\theta}-\sin{\theta}\right)}{16{n_{0}}^{2}}+$ (S17)
@@
$\langle a^{2}\rangle$ $=n_{0}\Big{(}1+\frac{\sin^{2}{\frac{\theta}{2}}}{{n_{0}}}+\frac{\left(1+2\bar{n}\right)[\left(1+\frac{\theta^{2}}{4}\right)\cos{\theta}-1]+\left(3+2\bar{n}\right)\frac{\theta}{4}\sin{\theta}-\frac{\theta^{2}}{2}}{4{n_{0}}^{2}}$ (S18)
@@
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+1-P_{e}\left(t\right)$ (S19)
Using the expansion of $P_{e}\left(t\right)$ up to second order Eq. (S7), we simplify the above term as
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+\sin^{2}{\frac{\theta}{2}}+\frac{\theta\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)}{16n_{0}}+\frac{\theta\left(1+\bar{n}\right)\left(1+2\bar{n}\right)\left(\theta\cos{\theta}-\sin{\theta}\right)}{16n_{0}^{2}}$ (S20)
@@
$V$ $=\begin{bmatrix}\sigma_{q}^{2}&\sigma_{qp}\\ \sigma_{pq}&\sigma_{p}^{2}\\ \end{bmatrix}$ (S21)
@@
[…]
@@
$V$ $=\begin{bmatrix}\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}-2\langle a\rangle^{2}&0\\ 0&-\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}\\ \end{bmatrix}\;.$ (S22)
For Gaussian states, the von Neumann entropy can be evaluated as
$S_{C}\left(t\right)$ $=h\left(\sqrt{Det[V]}\right),\hskip 5.0pth\left(x\right)=\left(x+\frac{1}{2}\right)\log\left({x+\frac{1}{2}}\right)-\left(x-\frac{1}{2}\right)\log\left({x-\frac{1}{2}}\right)\;.$ (S23)
@@
$\sqrt{Det[V]}$ $=\left(\bar{n}+\frac{1}{2}\right)\sqrt{[1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)]}$ (S24)
Using the framework of [30] we can calculate the heat exchanges of the cavity. We must first find the thermal state $\hat{\rho}_{C}^{th}\left(t\right)$ with the same von Neumann entropy $S_{C}\left(t\right)$ to do so. The von Neumann entropy of a thermal state with photon numbers $n_{c}$ is given by $\log[\frac{n_{c}^{n_{c}}}{\left(n_{c}+1\right)^{n_{c}+1}}]$. Comparing the entropies
$\log[\frac{n_{c}^{n_{c}}}{\left(n_{c}+1\right)^{n_{c}+1}}]=S_{C}\left(t\right)\implies n_{c}+\frac{1}{2}=\sqrt{Det[V]}$
To find the heat exchanges $Q_{C}$, we recall that the heat is defined as the negative of the variation in the thermal energy of the cavity, which takes the following form $Q_{C}\left(t\right)=-Tr[\hat{H}_{C}\left(\hat{\rho}_{C}^{th}\left(t\right)-\hat{\rho}_{C}^{th}\left(0\right)\right)]$ and reads
$Q_{C}\left(t\right)=\hbar\omega_{0}\left(\bar{n}+\frac{1}{2}\right)\left[1-\sqrt{1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)}\right]\;.$ (S25)
@@
$\Delta E_{C}\left(t\right)=\hbar\omega_{0}\left(1-P_{e}\left(t\right)\right)$ (S26)
@@
$W_{C}\left(t\right)=-\Delta E_{C}\left(t\right)-Q_{C}\left(t\right)\;.$ (S27)
III C. Joint atom-cavity evolution on longer times
III.1 Measurement by the cavity
In this section, we provide the expressions for cavity and atom, evolution over the interval $[0,t_{c}]$, when the system starts in state:
$\hat{\rho}\left(0\right)=\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{C}\left(0\right),\hskip 5.0pt\hat{\rho}_{C}\left(0\right)=D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}$ (S28)
@@ III.1 Measurement by the cavity
|±$|\pm_{y}\rangle\langle\pm_{y}|=\frac{1}{2}\begin{bmatrix}1&\mp i\\ \pm i&1\end{bmatrix},\hskip 10.0pt|\pm_{y}\rangle\langle\mp_{y}|=\frac{1}{2}\begin{bmatrix}1&\pm i\\ \pm i&-1\end{bmatrix}.$ (S29)
@@
$\hat{U}\left(t\right)=\hat{c}_{1}|e\rangle\langle e|+\hat{c}_{0}|g\rangle\langle g|+\hat{s}_{1}|e\rangle\langle g|-\hat{s}_{0}|g\rangle\langle e|,$ (S30)
@@
$\hat{s}_{1}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}},\hskip 5.0pt\hat{s}_{0}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (S31)
@@
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\hat{U}\left(t\right)\frac{1}{2}[|e\rangle\langle e|-i|e\rangle\langle g|+i|g\rangle\langle e|+|g\rangle\langle g|]\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ (S32)
$=\sum_{p,q=e,g}|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$
@@
where $\hat{T}_{pq}=|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{C}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$. For $p=e$ and $q=e$:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{1}^{\dagger}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{s}_{1}^{\dagger}+i\hat{s}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}].$ (S33)
@@
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{C}\left(0\right)a^{\dagger}\hat{s}_{1}^{{}^{\prime}}-i\hat{c}_{1}\hat{\rho}_{C}\left(0\right)a^{\dagger}\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{C}\left(0\right)\hat{c}_{1}]$ (S34)
$=\frac{1}{2}|e\rangle\langle e|\otimes\Big{[}\hat{c}_{1}\hat{\rho}_{C}\left(0\right)\hat{c}_{1}+$
$\hat{s}_{1}^{{}^{\prime}}[n_{0}\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)]\hat{s}_{1}^{{}^{\prime}}$
$-i\hat{c}_{1}\left[\alpha^{*}\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\right]\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}[\alpha\hat{\rho}_{C}\left(0\right)+D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)]\hat{c}_{1}\Big{]}$
The terms that include at least one bosonic operator acting on the thermal state, such as $D\left(\alpha_{0}\right)\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha$, $D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}$, and $D\left(\alpha_{0}\right)a\hat{w}_{\beta\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)$, have a vanishing magnitude when tracing over the cavity, and are therefore neglected, leading to:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\alpha\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{1}-i\alpha^{*}\hat{s}_{1}^{{}^{\prime}}\right)$ (S35)
$=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{C}\left(0\right)\left(\hat{c}_{1}-i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)$
@@
$\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}+1}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}}$ (S36)
@@
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{C}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right).$ (S37)

### v4

@@ Appendix B Quasi-ideal work source behavior of the cavity
$\hat{\rho}_{\text{C}}\left(t\right)$ $=\cos\left(gt\sqrt{a^{\dagger}a+1}\right)D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\cos\left(gt\sqrt{a^{\dagger}a+1}\right)+$ (39)
$\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}a^{\dagger}D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a\frac{\sin\left({gt\sqrt{a^{\dagger}a}}\right)}{\sqrt{a^{\dagger}a}}$
@@
$\langle a\rangle$ $=Tr[\hat{\rho}_{\text{C}}\left(t\right)a]$ (40)
$=\sum_{n}[\cos\left(gt\sqrt{n+1}\right)\cos\left(gt\sqrt{n}\right)+{\sin\left({gt\sqrt{n+1}}\right)}\sin\left({gt\sqrt{n}}\right)\frac{\sqrt{n+1}}{\sqrt{n}}]\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle$
@@
$\langle a\rangle$ $=\sum_{n}\left(\Lambda_{0}+\Lambda_{1}n+\Lambda_{2}n^{2}\right)\langle n|D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)a|n\rangle+O\left(\frac{1}{n_{0}^{4}}\right)$ (41)
@@
$\Lambda_{0}=\frac{{1+\frac{\theta^{2}}{4}+2n_{0}\left(-3+4n_{0}\left(3+4n_{0}\right)-\theta^{2}\right)+\left(-1+6n_{0}-24n_{0}^{2}+\left(1-2n_{0}\right)^{2}\frac{\theta^{2}}{4}\right)\cos\left(\theta\right)-\left(1-3n_{0}\right)^{2}\theta\sin\left(\theta\right)}}{{32n_{0}^{3}}}$ (42)
$\Lambda_{1}=\frac{{1-6n_{0}+\frac{\theta^{2}}{4}+\left(-1+\frac{\theta^{2}}{4}-2n_{0}\left(-3+\frac{\theta^{2}}{4}\right)\right)\cos\left(\theta\right)+\left(-2+7n_{0}\right)\frac{\theta}{2}\sin\left(\theta\right)}}{{8n_{0}^{3}}}$ (43)
$\Lambda_{2}=\frac{{4+2\left(\frac{\theta^{2}}{4}-2\right)\cos\left(\theta\right)-\frac{5\theta}{2}\sin\left(\theta\right)}}{{16n_{0}^{3}}}$ (44)
@@
$\langle a\rangle$ $=\sqrt{n_{0}}\Big(1+\frac{\sin^{2}{\frac{\theta}{2}}}{2{n_{0}}}+\frac{-2\sin^{2}{\frac{\theta}{2}}\left(1+\frac{\theta^{2}}{2}\right)+\frac{\theta}{2}\sin{\theta}+\bar{n}\theta\left(\theta\cos{\theta}-\sin{\theta}\right)}{16{n_{0}}^{2}}+$ (45)
@@
$\langle a^{2}\rangle$ $=n_{0}\Big(1+\frac{\sin^{2}{\frac{\theta}{2}}}{{n_{0}}}+\frac{\left(1+2\bar{n}\right)[\left(1+\frac{\theta^{2}}{4}\right)\cos{\theta}-1]+\left(3+2\bar{n}\right)\frac{\theta}{4}\sin{\theta}-\frac{\theta^{2}}{2}}{4{n_{0}}^{2}}$ (46)
@@
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+1-P_{e}\left(t\right)$ (47)
Using the expansion of $P_{e}\left(t\right)$ up to second order Eq. (33), we simplify the above term as
$\langle a^{\dagger}a\rangle$ $=\bar{n}+n_{0}+\sin^{2}{\frac{\theta}{2}}+\frac{\theta\left(\left(1+2\bar{n}\right)\theta\cos{\theta}+\left(3+2\bar{n}\right)\sin{\theta}\right)}{16n_{0}}+\frac{\theta\left(1+\bar{n}\right)\left(1+2\bar{n}\right)\left(\theta\cos{\theta}-\sin{\theta}\right)}{16n_{0}^{2}}$ (48)
@@
$V$ $=\begin{bmatrix}\sigma_{q}^{2}&\sigma_{qp}\\ \sigma_{pq}&\sigma_{p}^{2}\\ \end{bmatrix}$ (49)
@@
$\sigma_{xy}=\frac{1}{2}\langle\hat{x}\hat{y}+\hat{y}\hat{x}\rangle-\langle\hat{x}\rangle\langle\hat{y}\rangle,\hskip 5.0pt\left(x,y\right)\in\{q,p\},\hskip 1.0pt\hat{q}=\frac{a+a^{\dagger}}{\sqrt{2}},\hskip 1.0pt\hat{p}=\frac{a-a^{\dagger}}{i\sqrt{2}}$
@@
$V$ $=\begin{bmatrix}\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}-2\langle a\rangle^{2}&0\\ 0&-\langle a^{2}\rangle+\langle a^{\dagger}a\rangle+\frac{1}{2}\\ \end{bmatrix}\;$. (50)
For Gaussian states, the von Neumann entropy can be evaluated as [50]
$S_{\text{C}}\left(t\right)$ $=h\left(\sqrt{Det[V]}\right),\hskip 5.0pth\left(x\right)=\left(x+\frac{1}{2}\right)\log\left({x+\frac{1}{2}}\right)-\left(x-\frac{1}{2}\right)\log\left({x-\frac{1}{2}}\right)\;.$ (51)
@@
$\sqrt{Det[V]}$ $=\left(\bar{n}+\frac{1}{2}\right)\sqrt{[1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)]}$ (52)
Using the framework of [30] we can calculate the heat exchanges of the cavity. We must first find the thermal state $\hat{\rho}_{\text{C}}^{th}\left(t\right)$ with the same von Neumann entropy $S_{\text{C}}\left(t\right)$ to do so. The von Neumann entropy of a thermal state with photon numbers $n_{\text{C}}$ is given by $\log[\frac{n_{\text{C}}^{n_{\text{C}}}}{\left(n_{\text{C}}+1\right)^{n_{\text{C}}+1}}]$. Comparing the entropies
$\color[rgb]{0,0,0}{-\log[\frac{n_{\text{C}}^{n_{\text{C}}}}{\left(n_{\text{C}}+1\right)^{n_{\text{C}}+1}}]=S_{\text{C}}\left(t\right)\implies n_{\text{C}}+\frac{1}{2}=\sqrt{Det[V]}}$}
To find the heat exchanges $Q_{\text{C}}$, we recall that the heat is defined as the negative of the variation in the thermal energy of the cavity, which takes the following form $Q_{\text{C}}\left(t\right)=-Tr[\hat{H}_{\text{C}}\left(\hat{\rho}_{\text{C}}^{th}\left(t\right)-\hat{\rho}_{\text{C}}^{th}\left(0\right)\right)]$ and reads
$Q_{\text{C}}\left(t\right)=\hbar\omega_{0}\left(\bar{n}+\frac{1}{2}\right)\left[1-\sqrt{1+\frac{1}{4n_{0}\left(2\bar{n}+1\right)}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)}\right]$ (53)
which up to first order in $1/n_{0}$ yields:
$Q_{\text{C}}\left(t\right)=-\frac{\hbar\omega_{0}}{16n_{0}}\left(\theta^{2}+\sin^{2}{\theta}+2\left(1+2\bar{n}\right)\theta\sin{\theta}\right)+O(\frac{1}{n_{0}^{2}})\;.$ (54)
@@
$\Delta E_{\text{C}}\left(t\right)=\hbar\omega_{0}\left(1-P_{e}\left(t\right)\right)$ (55)
@@
$W_{\text{C}}\left(t\right)=-\Delta E_{\text{C}}\left(t\right)-Q_{\text{C}}\left(t\right)\;.$ (56)
Using Eq. (37) and Eq. (54), we can write the entropy production denoted as $\sigma_{\text{Q}}$ of the qubit system as
$\sigma_{\text{Q}}=S_{\text{Q}}(t)-\beta_{C}(0)Q_{\text{C}}(t)\geq 0$ (57)
Appendix C Joint atom-cavity evolution on longer times
C.1 Perturbative expansion of the global atom–cavity state
In this section, we provide the expressions for cavity and atom evolution over the interval $0,t_{c}]$, when the system starts in state:
$\hat{\rho}\left(0\right)=\sum_{\nu,\mu=\pm}\rho_{\nu,\mu}|\nu_{y}\rangle\langle\mu_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right),\hskip 5.0pt\hat{\rho}_{\text{C}}\left(0\right)=D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right),\hskip 5.0pt\alpha_{0}=\sqrt{n_{0}}$ (58)
@@ C.1 Perturbative expansion of the global atom–cavity state
$|\pm_{y}\rangle\langle\pm_{y}|=\frac{1}{2}\begin{bmatrix}1&\mp i\\ \pm i&1\end{bmatrix},\hskip 10.0pt|\pm_{y}\rangle\langle\mp_{y}|=\frac{1}{2}\begin{bmatrix}1&\pm i\\ \pm i&-1\end{bmatrix}$. (59)
@@
$\hat{U}\left(t\right)=\hat{c}_{1}|e\rangle\langle e|+\hat{c}_{0}|g\rangle\langle g|+\hat{s}_{1}|e\rangle\langle g|-\hat{s}_{0}|g\rangle\langle e|,$ (60)
@@
$\hat{s}_{1}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}},\hskip 5.0pt\hat{s}_{0}^{{}^{\prime}}=\frac{\sin\left(gt\sqrt{\hat{n}}\right)}{\sqrt{\hat{n}}}$ (61)
@@
$\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ $=\hat{U}\left(t\right)\frac{1}{2}[|e\rangle\langle e|-i|e\rangle\langle g|+i|g\rangle\langle e|+|g\rangle\langle g|]\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)$ (62)
$=\sum_{p,q=e,g}|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$
@@
where $\hat{T}_{pq}=|p\rangle\langle q|\otimes\langle p|\hat{U}\left(t\right)|+_{y}\rangle\langle+_{y}|\otimes\hat{\rho}_{\text{C}}\left(0\right)\hat{U}^{\dagger}\left(t\right)|q\rangle$. For $p=e$ and $q=e$:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{1}+\hat{s}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{s}_{1}^{\dagger}-i\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{s}_{1}^{\dagger}+i\hat{s}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{1}].$ (63)
@@
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes[\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{1}+\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{\text{C}}\left(0\right)a^{\dagger}\hat{s}_{1}^{{}^{\prime}}-i\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)a^{\dagger}\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}a\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{1}]$ (64)
$=\frac{1}{2}|e\rangle\langle e|\otimes\Big[\hat{c}_{1}\hat{\rho}_{\text{C}}\left(0\right)\hat{c}_{1}+$
$\hat{s}_{1}^{{}^{\prime}}[n_{0}\hat{\rho}_{\text{C}}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha+D\left(\alpha_{0}\right)a\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}+D\left(\alpha_{0}\right)a\hat{w}_{\beta_{C}\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)]\hat{s}_{1}^{{}^{\prime}}$
$-i\hat{c}_{1}\left[\alpha^{*}\hat{\rho}_{\text{C}}\left(0\right)+D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\right]\hat{s}_{1}{{}^{\prime}}+i\hat{s}_{1}^{{}^{\prime}}[\alpha\hat{\rho}_{\text{C}}\left(0\right)+D\left(\alpha_{0}\right)a\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)]\hat{c}_{1}\Big]$
The terms that include at least one bosonic operator acting on the thermal state, such as $D\left(\alpha_{0}\right)\hat{w}_{\beta_{C}\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)\alpha$, $D\left(\alpha_{0}\right)a\hat{w}_{\beta_{C}\left(0\right)}D^{\dagger}\left(\alpha_{0}\right)\alpha^{*}$, and $D\left(\alpha_{0}\right)a\hat{w}_{\beta_{C}\left(0\right)}a^{\dagger}D^{\dagger}\left(\alpha_{0}\right)$, have a vanishing magnitude when tracing over the cavity, and are therefore neglected, leading to:
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\alpha\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{\text{C}}\left(0\right)\left(\hat{c}_{1}-i\alpha^{*}\hat{s}_{1}^{{}^{\prime}}\right)$ (65)
$=\frac{1}{2}|e\rangle\langle e|\otimes\left(\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)\hat{\rho}_{\text{C}}\left(0\right)\left(\hat{c}_{1}-i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}\right)$
@@
$\hat{c}_{1}+i\sqrt{n_{0}}\hat{s}_{1}^{{}^{\prime}}$ $=\cos\left(gt\sqrt{\hat{n}+1}\right)+i\sqrt{n_{0}}\frac{\sin\left(gt\sqrt{\hat{n}+1}\right)}{\sqrt{\hat{n}+1}}$ (66)
@@
$\hat{T}_{ee}$ $=\frac{1}{2}|e\rangle\langle e|\otimes e^{\frac{igt\hat{n}}{2\sqrt{n_{0}}}}\hat{\rho}_{\text{C}}\left(0\right)e^{-\frac{igt\hat{n}}{2\sqrt{n_{0}}}}+O\left(\frac{1}{n_{0}}\right).$ (67)